# One Permutation Is All You Need: Fast, Reliable Variable Importance and Model Stress-Testing

### v1

@@ 3.2.2 VI Methods Considered
We test both variants of our proposed method: the one using the optimal rank-shift permutation, and the one using the simpler index-shift permutation, which acts as a cheaper approximation to the optimal permutation scheme. As baseline we consider the classical Breiman-style permutation importance methods implemented in standard Python libraries like scikit-learn, both using one random repetition (the fastest possible setup) and the much more common default of ten (Alonso2022). We use the same MSE-based scoring metric in all cases so that differences in results cannot be attributed to differences in scoring metrics. However, for completeness, we include in the Appendix the results obtained using default scoring metrics, which are even more favorable to our proposed methods.
@@ 5 From Variable Importance to Model Stress-Testing
In any case, it makes sense to consider a minimum correlation threshold beyond which propagate perturbations, as features may exhibit non-zero empirical correlation even if their true (population) correlation is zero. We propose estimating such threshold as a user-defined quantile (default: 0.9) of the absolute correlation scores among synthetic features randomly sampled from a multivariate standard Gaussian distribution with identity covariance matrix, which acts as a null distribution. Hence, the surviving correlations are likely not spurious – intuitively (although not formally) with confidence equal to the selected quantile. In order to make results deterministic, we internally fix a random seed that cannot be modified externally by the user.
@@ 5.1 Case Study A: Debt-to-Income Ratio Estimation
We randomized the order of the rows in the original dataset and performed 10-fold cross-validation to obtain repeated measures of master model accuracy as well as VI method accuracy.
@@ 5.1.1 Direct VI
Tables 15 and 16 in the Appendix show the results employing default scoring functions for our methods and the classical baseline (but in the case of the latter it is the same scoring function in the case of regression). Those results highlight that correlation with ground truth is rather stable across scoring functions, but that is not the case for a more extreme metric like the maximum score difference.
@@ 5.1.2 Systemic VI
  • With a plain OLS master, the systemic relative importance of feature black in the fitted model is estimated to be 1.46%, which can be decomposed as the sum of its direct (0.03%) and indirect (1.43%) relative importance scores. Thus, its systemic importance is over 40 times larger than its direct importance (which is what classical Breiman-style methods would only be able to estimate). Using MAE instead of MSE, we obtain the following importance decomposition: 4.80% = 0.92% + 3.88%, which in this case implies a smaller multiplier of 5.
  • With a sparse linear regressor master, the systemic relative importance of feature black in the fitted model is estimated to be 1.11%, which is due entirely to its indirect component, as the direct importance of the feature in the model is exactly 0. Using MAE instead of MSE, we obtain a larger systemic importance score of 4.19%, again entirely due to network effects.
Hence, even if the sensitive attribute black is not used directly by the model, our systemic importance method identified that the model still relies on it indirectly through correlated covariates beyond any spurious correlation doubts (tolerance quantile: 0.99), such as lvr (ratio of size of loan to assessed value of property), css (consumer credit score) or deny (1 if the applicant had a mortgage denied, 0 otherwise). Our method quantified this indirect reliance on the protected attribute to be between 1% and 4% of the total predictive reliance of the model, despite not being present in the model as such. This reliance on a protected attribute via proxy variables would be an undetected model risk and potential regulatory non-compliance when assessed using traditional VI methods.
@@
For completeness, we also tested fitting the sparse linear regressor master with all covariates and the results are similar to the OLS case. The profile comparison report made with the trust-free library shows that the direct counterfactual effect on prediction of black indeed typically is about 1% (see Figure 19 in the Appendix).
@@ 5.2 Case Study B: Credit Risk Forecasting
Again, being able to define ground-truth VI scores requires using a model class in which they are undisputedly identifiable, such as linear models (in log-odds). Hence we consider $\ell_{1}$-regularized logistic regression (with a strong penalty so that most coefficients are zeroed-out) and unregularized logistic regression (so that all feature coefficients are kept nonzero). We also consider RF as a black-box benchmark to illustrate one of the main use cases of VI methods. As in the previous regression case study, we confirm that our proposed methods achieve a similar or slightly higher correlation with the RF-native VI scores compared to the classical baseline methods (0.54-57 in all cases).
@@ 5.2.1 Direct VI
Tables 17 and 18 in the Appendix show the results employing default scoring functions for our methods and the classical baseline. In that case, results favor very significantly our proposed methods compared to baseline.
@@ 5.2.2 Systemic VI
  • With an unregularized logistic regression master, the systemic relative importance of feature Sex-Marital_status in the fitted model is estimated to be 1.37%, which can be decomposed as the sum of its direct (3.10%) and indirect (-1.73%) relative importance scores. Observe that in this case the direct importance is about twice as large as the systemic importance of the feature, indicating a decoupling effect: the protected attribute shows weaker-than-average interconnection with the rest of covariates, providing critical insight into the feature’s specific role within the network (which would not be accessible with classical Breiman-style VI methods). Using MAE instead of MSE, we obtain the following importance decomposition: 3.17% = 5.05% - 1.88%, which has similar qualitative interpretation as before.
  • With a sparse logistic regression master that excludes the protected feature, the systemic relative importance of said feature in the fitted model is estimated to be essentially non-existent at 0.01%. Using MAE instead of MSE, the estimated systemic importance is still 0 for all intents and purposes (0.36%), showing effective safeguarding by exclusion, in this case.
@@
Figure 2 shows the strength of interconnections among covariates, including the sensitive attribute Sex-Marital_status.
@@ 6 Conclusion
As machine learning models become increasingly central to high-stakes decision-making in finance and macroeconomic policy, the tools used to interpret and stress-test them must meet rigorous standards of stability, efficiency, and reliability. In this paper, we challenged the long-standing convention that permutation-based variable importance requires repeated random sampling to be effective.
@@ Appendix A Results with Default Scoring Metrics
Appendix B Additional Figures and Plots
@@
Appendix C Counterfactual Analysis With the trust-free Python Library

### v2

@@ 3.2.2 VI Methods Considered
We test both variants of our proposed method: the one using the optimal rank-shift permutation, and the one using the simpler index-shift permutation, which acts as a cheaper approximation to the optimal permutation scheme. As baseline we consider the classical Breiman-style permutation importance methods implemented in standard Python libraries like scikit-learn, both using one random repetition (the fastest possible setup) and the much more common default of ten (Alonso2022). We use the same MSE-based scoring metric in all cases so that differences in results cannot be attributed to differences in scoring metrics. However, for completeness, we include in Appendix A the results obtained using default scoring metrics, which are even more favorable to our proposed methods.
@@ 5 From Variable Importance to Model Stress-Testing
5.1 Statistical Calibration of the Propagation Threshold
To ensure that systemic variable importance (SVI) captures structural dependencies rather than numerical artifacts, we implement a non-parametric calibration mechanism to establish a minimum correlation threshold, $\tau$. In any finite dataset, features may exhibit non-zero empirical correlations due to sampling error, even when their true population-level associations are zero.
5.1.1 Hypothesis Testing
We define the global null hypothesis of independence, $H_{0}$, such that the joint distribution of any two features $X_{i}$ and $X_{j}$ factors into the product of their marginals: $f(x_{i},x_{j})=f(x_{i})f(x_{j})$. To maintain the integrity of the forensic audit, the propagation of perturbations is restricted to feature pairs whose association is statistically significant relative to the permutation distribution. We seek to learn a threshold $\tau$ that provides a family-wise safeguard, ensuring that the probability of any spurious correlations exceeding this limit is controlled at a global significance level $\alpha$:
$\mathbb{P}\left(\max_{1\leq i<j\leq p}|\hat{\rho}_{ij}|>\tau\mid H_{0}\right)\leq\alpha$ (19)
5.1.2 Non-parametric Estimation
In practice, we estimate $\tau$ in-sample using an efficient permutation-based approach. Let $\mathbf{X}\in\mathbb{R}^{n\times p}$ be the training matrix. We generate a null representation by independently permuting the row indices of each column, creating a synthetic dataset $\tilde{\mathbf{X}}$. This procedure preserves the marginal distributions $f(x_{k})$ for all $k\in\{1,\dots,p\}$ while stochastically breaking all inter-feature dependencies.
Under $H_{0}$, the M=p(p1)/2M=p(p-1)/2 pairwise correlations in the upper triangular matrix are exchangeable. We therefore define $\tau$ as the $(1-\alpha)$ quantile of this pooled empirical distribution:
$\tau=Q\left(\{|\hat{\rho}(\tilde{X}_{i},\tilde{X}_{j})|\}_{1\leq i<j\leq p}\,;\,1-\alpha\right)$ (20)
where $Q(S;p)$ is the empirical quantile function defined as the infimum of the set of values whose empirical cumulative distribution $F_{S}$ is at least $p$:
$Q(S;p)=\inf\{x\in\mathbb{R}:F_{S}(x)\geq p\}$ (21)
Note that utilizing the inverted cumulative distribution function ensures that $\tau$ is an observed value from the permutation set, which provides a conservative family-wise safeguard that strictly respects the empirical noise floor.
By pooling all pairwise entries from the permuted matrix, $\tau$ provides a robust estimate of the global noise floor while accounting for the multiple comparisons inherent in auditing the full feature space. This approach provides a computationally efficient estimator for strong Family-Wise Error Rate (FWER) control (i.e., control under any configuration of the true dependencies). Correlations exceeding $\tau$ are treated as systemic proxies, whereas those below $\tau$ are dismissed as sampling artifacts. To ensure the audit remains deterministic and reproducible, the internal permutation seed is fixed.
@@ 5.2 Case Study A: Debt-to-Income Ratio Estimation
We randomize the order of the rows in the original dataset and perform 10-fold cross-validation to obtain repeated measures of master model accuracy as well as direct VI method accuracy.
For consistency with our 10-fold cross-validation, we learn the spurious correlation threshold using the training set while we estimate SVI on the 10% held-out test set.
@@ 5.2.1 Direct VI
Tables 15 and 16 in Appendix A show the results employing default scoring functions for our methods and the classical baseline (but in the case of the latter it is the same scoring function in the case of regression). Those results highlight that correlation with ground truth is rather stable across scoring functions, but that is not the case for a more extreme metric like the maximum score difference.
Lastly, to further test the stability of the traditional VI methods on real datasets we execute them ten times and check the top 5 features that they report each time. Table 19 in Appendix B shows that the classical baseline employing only one repetition yields very unstable rankings for the top 5 features: every single run yields a different ranking. Employing 10 repetitions reduces this instability to some degree although it is still notable: 7 distinct rankings were produced in 10 runs. In stark contrast, our deterministic framework is designed to yield the exact same scores (and, hence, rankings) every single time it is run. In addition, only our proposed methods (either with optimal or sub-optimal permutation) recovered the ground-truth top 5 components.
@@ 5.2.2 Systemic VI
  • With a plain OLS master, the systemic relative importance of feature black in the fitted model is estimated to be 0.44%, which can be decomposed as the sum of its direct (0.09%) and indirect (0.35%) relative importance scores. Thus, its systemic importance is about 5 times larger than its direct importance (which is what classical Breiman-style methods would only be able to estimate). Using MAE instead of MSE, we obtain the following importance decomposition: 2.31% = 1.21% + 1.10%, which in this case implies a smaller multiplier of 2.
  • With a sparse linear regressor master, the systemic relative importance of feature black in the fitted model is estimated to be 0.16%, which is due entirely to its indirect component, as the direct importance of the feature in the model is exactly 0. Using MAE instead of MSE, we obtain a larger systemic importance score of 1.41%, again entirely due to network effects.
Hence, even if the sensitive attribute black is not used directly by the model, our systemic importance method identified that the model still relies on it indirectly through correlated covariates beyond any spurious correlation doubts (tolerance quantile: 0.99), such as lvr (ratio of size of loan to assessed value of property), css (consumer credit score) or deny (1 if the applicant had a mortgage denied, 0 otherwise). Our method quantified this indirect reliance on the protected attribute to be between 0.16% and 1.41% of the total predictive reliance of the model, despite not being present in the model as such. This reliance on a protected attribute via proxy variables would be an undetected model risk and potential regulatory non-compliance when assessed using traditional VI methods.
@@
Observe that the protected attribute black shows Spearman correlations of about 0.2 or higher with 4 of the 12 features included in the dataset: deny (whether the mortgage application was denied), css (consumer credit score from 1 to 6, where a lower value represents a better score), lvr (loan-to-value ratio), and condominium (whether the applicant lives in a condominium as opposed to a detached single-family home). This means that the model can “reconstruct” a sizable proportion of the racial information using those proxies, even when the race variable is explicitly excluded from the model.
More generally, taking into account all features having Spearman correlation with black exceeding the 0.99 quantile from a non-parametric null distribution to define its significant proxies, we note that 90% of the total ground-truth score contributions in a model that excludes the race variable is nonetheless influenced by race.
For completeness, we also tested fitting the sparse linear regression master with all covariates and the results are similar to the OLS case. The profile comparison report made with the trust-free library shows that the direct counterfactual effect on prediction of black indeed typically is about 1% (see Figure 19 in Appendix D).
@@ 5.3 Case Study B: Credit Risk Forecasting
Again, being able to define ground-truth VI scores requires using a model class in which they are undisputedly identifiable, such as linear models (in log-odds). Hence we consider $\ell_{1}$-regularized logistic regression (with a strong penalty so that most coefficients are zeroed-out) and unregularized logistic regression (so that all feature coefficients are kept nonzero). We also consider RF as a black-box benchmark to illustrate one of the main use cases of VI methods. As in the previous regression case study, we confirm that our proposed methods achieve a similar correlation with the RF-native VI scores compared to the classical baseline methods (0.50-52 in all cases).
@@ 5.3.1 Direct VI
Tables 17 and 18 in Appendix A show the results employing default scoring functions for our methods and the classical baseline. In that case, results favor very significantly our proposed methods compared to baseline.
@@
Similarly to the previous case study, Table 20 in Appendix B shows that across 10 independent runs on the German Credit dataset, both our optimal and approximate VI methods recovered the exact ground-truth top-5 feature set in all runs. On the other hand, both classical baselines employing one or even ten repetitions yielded very unstable rankings for the top 5 features, producing 9 and 10 distinct top 5 rankings, respectively.
@@ 5.3.2 Systemic VI
  • With an unregularized logistic regression master, the systemic relative importance of feature Sex-Marital_status in the fitted model is estimated to be 3.24%, which can be decomposed as the sum of its direct (2.09%) and indirect (1.15%) relative importance scores. Thus, its systemic importance is about 50% larger than its direct importance. Using MAE instead of MSE, we obtain the following importance decomposition: 4.94% = 4.20% + 0.74%.
  • With a sparse logistic regression master that excludes the protected feature, the systemic relative importance of said feature in the fitted model is estimated to be exactly 0 as well. Using MAE instead of MSE, the estimated systemic importance is still 0, showing effective safeguarding by exclusion, in this case.
@@ 6 Conclusion
As machine learning models become increasingly central to high-stakes decision-making in finance and macroeconomic policy, the tools used to interpret and stress-test them must meet rigorous standards of stability, efficiency, and reliability. In this paper, we challenged the long-standing convention that permutation-based VI requires repeated random sampling to be effective.
@@ Appendix A Results with Default Scoring Metrics
Appendix B Stability Across Repeated Independent Runs
Table 19: Top 5 features by VI method, across 10 independent runs. Ground truth: $[0,1,11,6,3]$ – Boston HMDA dataset.
Top 5 indices Direct-Opt Direct-Approx Breiman (B=1) Breiman (B=10)
$[0,11,3,1,6]$ 10 10 0 0
$[0,11,2,9,3]$ 0 0 0 3
$[0,11,3,9,2]$ 0 0 0 2
$[0,11,9,2,1]$ 0 0 0 1
$[0,11,9,3,2]$ 0 0 0 1
$[0,11,1,2,9]$ 0 0 1 0
$[0,11,2,9,1]$ 0 0 1 0
$[0,11,1,2,7]$ 0 0 1 0
$[0,11,2,3,9]$ 0 0 1 0
$[0,3,11,9,1]$ 0 0 0 1
$[0,3,11,1,9]$ 0 0 0 1
$[0,2,3,9,11]$ 0 0 0 1
$[0,1,11,9,3]$ 0 0 1 0
$[0,3,1,9,7]$ 0 0 1 0
$[0,3,11,9,2]$ 0 0 1 0
$[0,3,11,9,7]$ 0 0 1 0
$[0,1,11,4,7]$ 0 0 1 0
$[0,3,2,11,9]$ 0 0 1 0
Table 20: Top 5 features across 10 independent runs. Ground truth: $[0,3,2,1,5]$ – German credit dataset.
Top 5 indices Direct-Opt Direct-Approx Breiman (B=1) Breiman (B=10)
$[0,3,2,1,5]$ 10 10 0 0
$[0,5,7,3,2]$ 0 0 2 0
$[0,19,5,18,3]$ 0 0 1 0
$[0,2,3,5,13]$ 0 0 1 0
$[0,7,5,15,18]$ 0 0 1 0
$[0,5,3,2,19]$ 0 0 1 0
$[0,7,9,4,18]$ 0 0 1 0
$[0,3,7,13,18]$ 0 0 0 1
$[0,3,7,2,18]$ 0 0 0 1
$[0,2,7,3,13]$ 0 0 0 1
$[0,2,7,5,3]$ 0 0 0 1
$[0,2,13,3,7]$ 0 0 0 1
$[0,7,5,3,13]$ 0 0 0 1
$[0,7,5,2,13]$ 0 0 0 1
$[0,7,2,5,3]$ 0 0 0 1
$[0,7,3,2,5]$ 0 0 0 1
$[0,5,7,2,13]$ 0 0 0 1
$[3,2,0,5,12]$ 0 0 1 0
$[3,13,19,0,18]$ 0 0 1 0
$[3,0,19,13,7]$ 0 0 1 0
Appendix C Additional Figures and Plots
@@
Appendix D Counterfactual Analysis With the trust-free Python Library